# Soft-QMIX: Integrating Maximum Entropy For Monotonic Value Function Factorization

### v2

@@ Appendix H Comparison with Baselines
2. Actor-critic methods utilize the KL divergence as the loss function, while our approach uses the MSE loss. In the CTDE framework, all methods maintain a more accurate centralized Q-function alongside multiple imprecise local policies. The core idea is to distill knowledge from the centralized Q-function to the local policies. Actor-critic methods typically minimize the KL divergence between the actor and the softmax critic, whereas our method minimizes the MSE loss between their logits. [12] pointed out that, in distillation contexts, MSE is a superior loss function compared to KL divergence.
@@
4. In fact, it’s challenging to theoretically prove that value-based methods are inherently superior to AC methods. However, through experiments, we have demonstrated that our method surpasses both FOP [36] and HASAC [16].
@@
mSAC [22] introduces a method similar to MASAC. It employs the AC architecture, thus encountering issues 1 and 2 mentioned above, but does not fully utilize global information to train the critic, hence lacking the advantage noted in point 3.
FOP [36] shares a similar issue with [22], in that the AC framework it utilizes can actually train the critic with global state information. Furthermore, FOP employs a credit assignment mechanism similar to that of QPLEX [31]. [9] experimentally shown that QMIX outperforms QPLEX. Additionally, FOP’s proof concept is that under the IGO conditions, if we locally move local policies towards local optimal policies, the distance between the joint policy and the optimal joint policy also decreases. In contrast, our proof approach first establishes that our method is equivalent to the single-agent SAC algorithm and then completes the proof using existing conclusions.
HASAC [16] is a CTCE version of MASAC, with its core contribution being the introduction of sequential decision-making. Therefore, we do not compare it here.
PAC [38] utilizes global information during centralized training, possessing the advantages noted in point 3. However, it still encounter issues 1 and 2 mentioned above.
Appendix I MPE setup
In the MPE [18] Spread scenario, three agents and three landmarks are present. Agents work together to cover all landmarks while avoiding collisions. We created three different maps: Spread(Simple), which matches the original benchmark’s simple_spread; Spread(Medium) and Spread(Hard), which increase the initial distances between agents and landmarks by 1.5X and 2X, respectively.
Appendix J Hyper-parameters
For all baseline algorithms, we implemented them using the corresponding open-source frameworks and chose the hyperparameters provided by these frameworks for replication. We used MAPPO implemented by  Huang et al., QMIX by  Hu et al., FOP by  Zhang et al., and HARL by  Zhong et al.. Our method, Soft-QMIX, shares the same code base as QMIX. The experiments were conducted on a computer equipped with 92 GB of RAM, a 40-core CPU, and a GeForce RTX 2080 Ti GPU. The following are the hyperparameters used in the experiments.
Hyperparameter Value
Learning Rate Actor 5e-4
Learning Rate Critic 1e-3
Data Chunk Length 8
Env Num 8
Episode Length 400
PPO Epoch 5
Actor Train Interval Step 1
Use Recurrent Policy True
Use Adv Normalize True
Use Value Active Masks False
Use Linear LR Decay True
Table 6: MAPPO hyper-parameters used for SMACv2. We utilize the hyperparameters used in SMACv2 [3].
@@ Appendix I MPE setup
Table 7: HASAC hyper-parameters used for SMACv2. We use the hyperparameters for SMAC as specified in the original paper [16].
@@
Action Selector epsilon greedy
$\epsilon$-Start 1.0
$\epsilon$-Finish 0.05
$\epsilon$-Anneal Time 1e5 (5e5 for MPE)
@@
Hypernet Embed Dimension 64
@@
Table 8: QMIX hyper-parameters used for SMACv2 and MPE. We utilize the hyperparameters used in SMACv2 [3].
@@
$\alpha$($\_5$ vs _ 5) 0.03
$\alpha$($\_10$ vs _ 10) 0.005
Table 9: Soft-QMIX hyper-parameters used for SMACv2. Note that besides the hyperparameters mentioned in the table, all other values for Soft-QMIX are the same as those for QMIX.
@@
Table 10: FOP hyper-parameters used for SMACv2. We use the hyperparameters for SMAC as specified in the original paper [36].
Appendix K Impact Statements
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.

### v3

@@ Acknowledgment
2. Actor-critic methods utilize the KL divergence as the loss function, while our approach uses the MSE loss. In the CTDE framework, all methods maintain a more accurate centralized Q-function alongside multiple imprecise local policies. The core idea is to distill knowledge from the centralized Q-function to the local policies. Actor-critic methods typically minimize the KL divergence between the actor and the softmax critic, whereas our method minimizes the MSE loss between their logits. (Kim et al. 2021) pointed out that, in distillation contexts, MSE is a superior loss function compared to KL divergence.
@@
4. In fact, it’s challenging to theoretically prove that value-based methods are inherently superior to AC methods. However, through experiments, we have demonstrated that our method surpasses both FOP (Zhang et al. 2021) and HASAC (Liu et al. 2023).
@@
mSAC (Pu et al. 2021) introduces a method similar to MASAC. It employs the AC architecture, thus encountering issues 1 and 2 mentioned above, but does not fully utilize global information to train the critic, hence lacking the advantage noted in point 3.
FOP (Zhang et al. 2021) shares a similar issue with (Pu et al. 2021), in that the AC framework it utilizes can actually train the critic with global state information. Furthermore, FOP employs a credit assignment mechanism similar to that of QPLEX (Wang et al. 2020). (Hu et al. 2021) experimentally shown that QMIX outperforms QPLEX. Additionally, FOP’s proof concept is that under the IGO conditions, if we locally move local policies towards local optimal policies, the distance between the joint policy and the optimal joint policy also decreases. In contrast, our proof approach first establishes that our method is equivalent to the single-agent SAC algorithm and then completes the proof using existing conclusions.
HASAC (Liu et al. 2023), in contrast to FOP (Zhang et al. 2021), avoids the restrictive Individual-Global-Optimal (IGO) assumption used for joint policy factorization by employing a theoretically grounded sequential optimization of individual policies via joint soft policy decomposition. However, our primary contribution addresses the misalignment problem, which is orthogonal to HASAC’s focus. Moreover, since HASAC does not rely on either the IGM or IGO assumptions, establishing a direct theoretical comparison is nontrivial. Nevertheless, we empirically demonstrate that our method consistently outperforms HASAC across benchmark tasks presented in the paper.
PAC (Zhou, Lan, and Aggarwal 2022) utilizes global information during centralized training, possessing the advantages noted in point 3. However, it still encounter issues 1 and 2 mentioned above.
Appendix E E. Hyper-parameters
For all baseline algorithms, we implemented them using the corresponding open-source frameworks and chose the hyperparameters provided by these frameworks for replication. We used FOP by  Zhang et al. and HASAC by Zhong et al.. Our method, ME-IGM is built on Hu et al.. The experiments were conducted on a computer equipped with 92 GB of RAM, a 40-core CPU, and a GeForce RTX 2080 Ti GPU. The following are the hyperparameters used in the experiments.
@@
Table 6: HASAC hyperparameters used for SMACv2. We use the hyperparameters for SMAC as specified in the original paper (Liu et al. 2023).
@@
Hyper-net Embed Dimension 64
@@
$\bar{H}$ (zerg, protoss) 0.24 $\times$ num_ally
$\bar{H}$ (terran) 0.32 $\times$ num_ally
$\alpha$ Learning Rate 0.3
Table 7: ME-QMIX hyperparameters used for SMACv2. We utilize the hyperparameters used in SMACv2 (Ellis et al. 2022).
@@
Runner parallel
Batch Size Run 8
Buffer Size 5000
Batch Size 128
Optimizer Adam
$t_{\max}$ 1005000
Target Update Interval 200
Mac basic_mac
Agent rnn
Agent Output Type q
Learner dmaq_qatten_learner
Mixer dmaq
Mixing Embed Dimension 32
Hyper-net Embed Dimension 64
Learning Rate 0.001
$\lambda$ 0.6
$\bar{H}$ (zerg, protoss) 0.24 $\times$ num_ally
$\bar{H}$ (terran) 0.32 $\times$ num_ally
$\alpha$ Learning Rate 0.3
Table 8: ME-QPLEX hyperparameters used for SMACv2. We utilize the hyperparameters used in SMACv2 (Ellis et al. 2022).
@@
Table 9: FOP hyper-parameters used for SMACv2. We use the hyperparameters for SMAC as specified in the original paper (Zhang et al. 2021).
Parameter Name Value
Runner parallel
Batch Size Run 4
Buffer Size 5000
Batch Size 512
Optimizer Adam
$t_{\max}$ 1005000
Target Update Interval 200
Mac n_mac
Agent n_rnn
Agent Output Type q
Learner nq_learner
Mixer qmix
Mixing Embed Dimension 32
Hyper-net Embed Dimension 64
Learning Rate 0.001
$\lambda$ 0.6
$\alpha$ Learning Rate 0.3
Table 10: ME-QMIX hyperparameters used for Overcooked. We utilize the hyperparameters used in Overcooked (Carroll et al. 2019).
Parameter Name Value
Runner parallel
Batch Size Run 8
Buffer Size 5000
Batch Size 512
Optimizer Adam
$t_{\max}$ 1005000
Target Update Interval 200
Mac basic_mac
Agent rnn
Agent Output Type q
Learner dmaq_qatten_learner
Mixer dmaq
Mixing Embed Dimension 32
Hyper-net Embed Dimension 64
Learning Rate 0.001
$\lambda$ 0.6
$\alpha$ Learning Rate 0.3
Table 11: ME-QPLEX hyperparameters used for Overcooked. We utilize the hyperparameters used in Overcooked (Carroll et al. 2019).